Suppose that in a certain neighborhood, the cost of a home (in thousands) is proportional to the size of the home in square feet. The regression equation quantifying this relationship is found to be (price) = 0.048*(size) + 39.137. You look more closely at one of the houses selected. The house is listed as having 1383.117 square feet and is listed at a price of $107.544 (thousand). The residual is 2.017. Interpret this residual in terms of the problem.

Question 5 options:

1)

The square footage is 2.017 square feet less than what we would expect.

2)

The price of the house is 2.017 thousand dollars larger than what we would expect.

3)

The price of the house is 2.017 thousand dollars less than what we would expect.

4)

The price of the house is 107.544 thousand dollars larger than what we would expect.

5)

The square footage is 2.017 square feet larger than what we would expect.
Answer #1

(price) = 0.048*(size) + 39.137

predicted price for size = 1383.117
(price) = 0.048*1383.117 + 39.137 = 105.53

actual price = 107.544


2) The price of the house is 2.017 thousand dollars larger than what we would expect.